1 mol C6H14 (hexane) is mixed with 1 mol C7H16 (heptane) at 298 K. p(C6H14)*=20.16 kPa, p(C7H16)*=6.25 kPa. Calculate the partial pressures and mole fractions of C6H14 and C7H16 in the gaseous phase separately at 298 K. Assume the solution and gas are ideal.
Answer #1

total pressure =  p(C6H14) + p(C7H16) = (20.16 + 6.25) = 26.41 KPa

and

mole fractions of C6H14 = p(C6H14) / total pressure = (20.16 / 26.41) = 0.763

and

mole fractions of C7H16 = (1 - 0.763) = 0.237